OpenStudy (anonymous):

A line has a slope of -1/2 and a y-intercept of 3. Which equations represent the line?

OpenStudy (anonymous):

A. 2y=6+x B. x+2y=6 C. y=-1/2x+3 D. (y-2)=-1/2(x-2)

OpenStudy (anonymous):

Choose all answers that are correct.

OpenStudy (anonymous):

\(\huge\rm\color{red}{y=mx+b}\) m=slope b=y intercept We know our slope is -1/2 and our y intercept (b) is 3 so what will the answer be once you sub in the values? :o

OpenStudy (anonymous):

C?

OpenStudy (anonymous):

A. 2y=6+x B. x+2y=6 D. (y-2)=-1/2(x-2) Good C is one of them now do you know how to write this equation in standered form?

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

Nope thats point slope form we dont have 2 points so it cant be that

OpenStudy (anonymous):

We are left with A and B

OpenStudy (anonymous):

A?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

That leaves us with B and C as our final answers.

OpenStudy (anonymous):

Thank You! ^.^ @Moe95

OpenStudy (anonymous):

No problem! :)

OpenStudy (anonymous):

Can you help with a few more please? @Moe95

OpenStudy (anonymous):

yes :)